OpenStudy (anonymous):

find the distance from (-6,-6) to the line defined by y=-2x-3

OpenStudy (anonymous):

Herethe y-intercept of the line y=-2x-3 is -3 therfore it is (0,-3) \[\sqrt{(-6-0)^{2 } +(-6+3)^{2}}\] \[\sqrt{36+9}=\sqrt{45}=3\sqrt{5}\]